OpenStudy (unklerhaukus):

Express as a percentage \[1°\textrm C / 100°\textrm C\]

hartnn (hartnn):

not 1% ?

OpenStudy (unklerhaukus):

not even close

OpenStudy (anonymous):

Have to convert it to something, haha?

OpenStudy (anonymous):

73%

OpenStudy (anonymous):

In K?

hartnn (hartnn):

units are same, its a ratio = 0.01 1% -_-

OpenStudy (unklerhaukus):

73% \(\checkmark\)

OpenStudy (anonymous):

No... 1°C = 273K

OpenStudy (anonymous):

1 degree Celsius = 273 Kelvin

OpenStudy (anonymous):

\[T = T(degree)+273.15 \implies 1+273.15 \] and \[100+273.15\] in kelvins divide and multiply by 100 73%

OpenStudy (anonymous):

K is the SI for temperature, iirc

OpenStudy (unklerhaukus):

celsius / centigrade is a terrible unit

OpenStudy (anonymous):

Haha, I agree, cool question UR!

hartnn (hartnn):

temp rised from 100 C to 101C I would say it increased by 1% not 73%

OpenStudy (anonymous):

Depends on the context. Math questions, ok. Physics questions, may not be.

OpenStudy (unklerhaukus):

\[101°\textrm C/100°\textrm C = 100.3\%\]

hartnn (hartnn):

in physics, rise in temp from 100C to 101C is 73% increase ?

hartnn (hartnn):

rise in temp from 100 C to 101C = (101-100)/100 = 1% increase

OpenStudy (anonymous):

\[\frac{(101+273.15)-(100+273.15)}{100+273.15}\]

OpenStudy (anonymous):

^Percentage change

hartnn (hartnn):

my whole graduation was a lie :O

hartnn (hartnn):

and post grad too :P